Electrochemical Measurement of Concentration An electrochemical cell consists of a standard hydrogen electrode and a copper mPlease show all work step by step and final answer.
Answer #1

The reduction of Cu2+ to Cu(s ) can be written as

           Cu2+(aq) + 2e-     ----> Cu(s)

A. The cell potential of the cell at 25 degree Celsius at different condition is given by Nernst equation

E = Eo- (0.0591/2)* log(1/[Cu2+])

= 0.34 - (0.0591/2)*log(1/3.6*10^-4) = 0.238 V

The cell potential of the cell at 25 degree Celsius with Cu2+ concentration of 3.6*10^-4 is 0.238 V

B. The cell potential of the cell at 25 degree Celsius at different condition is given by Nernst equation

E = Eo- (0.0591/2)* log(1/[Cu2+])

0.136 = 0.34 - (0.0591/2)*log(1/[Cu2+])

The concentration of the solution of unknown [Cu2+] = 1.25*10^-7 M

C. The cell potential of the cell at 25 degree Celsius at different condition is given by Nernst equation

E = Eo - (0.0591/2)* log(1/[Cu2+])

E = Eo + (0.0591/2)* log([Cu2+])


The above equation is equation of straight line(y = mx+c) with slope = 0.0591/2 and intercept = E0